# VisionUnite: A Vision-Language Foundation Model for Ophthalmology Enhanced with Clinical Knowledge

### v1

@@ Acknowledgements
Figure 6: The consistent interpretation analysis of visual and language features between large vision-language models and the junior ophthalmologist (Doctor in the figure). Red represents inaccurate or irrelevant descriptions. Blue represents the accurate descriptions. The expert (senior ophthalmologist) evaluation is also provided.
Figure 7: [Extended Data Figure 1] The joint diagnosis analysis of multiple Diseases or Conditions and signs. The figure shows the 95% Confidence Intervals (CI) for different models and the junior ophthalmologist (Doctor in the figure). All p-values are calculated with the two-sided T-test between VisionUnite and others. * means p-value < 0.1 and ** means p-value < 0.001.
Figure 8: [Extended Data Figure 2] The misdiagnosis analysis in healthy conditions. The top is the comparison of diagnostic accuracy, and the bottom is the comparison of diagnostic relevance. The figure shows the 95% Confidence Intervals (CI) for different models and the junior ophthalmologist (Doctor in the figure). All p-values are calculated with the two-sided T-test between VisionUnite and others. * means p-value < 0.1 and ** means p-value < 0.001.
Figure 9: [Extended Data Figure 3] The diagnostic correction analysis in patient interaction. Overall refers to answering correctly in the second or third round. Round 2 refers to answering correctly in the second round, while Round 3 refers to answering correctly in the third round. The figure shows the 95% Confidence Intervals (CI) for different models. All p-values are calculated with the two-sided T-test between VisionUnite and others. * means p-value < 0.1 and ** means p-value < 0.001.
Figure 10: [Extended Data Figure 4] The diagnostic continuity analysis in patient interaction includes incorrect prediction and correct prediction. The top is the comparison of diagnostic relevance with incorrect prediction, and the bottom is the comparison of diagnostic relevance with correct prediction. The figure shows the 95% Confidence Intervals (CI) for different models and the junior ophthalmologist (Doctor in the figure). All p-values are calculated with the two-sided T-test between VisionUnite and others. * means p-value < 0.1 and ** means p-value < 0.001.
Figure 11: [Extended Data Figure 5] The multiple-choice evaluation between LLaVA-Med and VisionUnite. The evaluation spans ten different diseases: Age-Related Macular Degeneration (AMD), Cataract, Central Serous Retinopathy (CSR), Diabetic Retinopathy (DR), Glaucoma, Media Haze, Myopia, Retinitis, Macular Edema and Tessellation. The numerical values indicate the multiple-choice accuracy of LLaVA-Med and VisionUnite on specific diseases.
Figure 12: [Extended Data Figure 6] The consistent interpretation analysis of case c and case d with the single Disease or Condition (age-related macular degeneration and diabetic retinopathy) between large vision-language models and the junior ophthalmologist (Doctor in the figure). Red represents inaccurate or irrelevant descriptions. Blue represents the accurate descriptions.
Figure 13: [Extended Data Figure 7] The consistent interpretation analysis of case e and case f with the single Disease or Condition (optic disk cupping and glaucoma) between large vision-language models and the junior ophthalmologist (Doctor in the figure). Red represents inaccurate or irrelevant descriptions. Blue represents the accurate descriptions.
Figure 14: [Extended Data Figure 8] The consistent interpretation analysis of case g and case h with multiple Diseases or Conditions (diabetic retinopathy/macular edema and diabetic retinopathy/optic disc pallor) between large vision-language models and the junior ophthalmologist (Doctor in the figure). Red represents inaccurate or irrelevant descriptions. Blue represents the accurate descriptions.
Figure 15: [Extended Data Figure 9] The consistent interpretation analysis of case i and case j with multiple Diseases or Conditions (diabetic retinopathy/arteriosclerotic retinopathy and myopia/optic disk cupping) between large vision-language models and the junior ophthalmologist (Doctor in the figure). Red represents inaccurate or irrelevant descriptions. Blue represents the accurate descriptions.
Figure 16: [Extended Data Figure 10] The corresponding expert (senior ophthalmologists) evaluation of case c-j, which includes the evaluation of diagnostic predictions and their relevance. $\times$ means incorrect prediction. $\checkmark$ means correct prediction. The expert assesses the relevance of each response set for diagnosis by ranking them from 4 to 1, based on their alignment with the label. A ranking of 4 indicates the highest consistency with the label, while a ranking of 1 indicates the lowest.
Appendix
      Private dataset
      (n = 33029)
      Variables       N       %
      Gender
      Female       17735       53.70%
      Male       15205       46.03%
      N/A       89       0.27%
      Age (years), Median (IQR)       39.75 (28.0-49.0)
      <60       29187       88.37%
      >=60       3014       9.12%
      N/A       828       2.51%
      Eye
      OD       16845       51.00%
      OS       16184       49.00%
      High myopia
      With       19592       59.32%
      Without       13437       40.68%
      Glaucoma
      With       5747       17.40%
      Suspect       4859       14.71%
      Without       22423       67.89%
      High intraocular pressure
      With       372       1.12%
      Suspect       12       0.04%
      Without       32645       98.84%
      Cup-to-Disc ratio increase
      With       1379       4.18%
      Without       31650       95.82%
      Cup-to-Disc ratio asymmetry
      With       237       0.72%
      Without       32792       99.28%
      Family history
      With       2272       6.88%
      Without       30757       93.12%
Table 1: [Table S1] Characteristics of patients in the Private dataset.
Table 2: [Table S2] The classification criteria for diagnostic errors include missing and incorrect errors, as well as minor and major errors. The classification criteria are applied to the analysis of diagnostic errors in this study.
The characteristics of patients in the private dataset: The detailed information of patients in the private dataset is shown in Table 1. The gender ratio and OD/OS ratio are balanced in the dataset. It also includes data on elderly patients aged 60 and above occupying 9.12% of the dataset.
Construction of the dialogue in the pretrain dataset
For the question section of the pretrain dataset, we have constructed 20 questions, categorized into two types: indicative long answers (10 sentences) and short answers (10 sentences). Depending on the length of the dialogue answer, a sentence from the relevant question type is selected as the interrogative component of the dialogue. All of the 20 constructed questions are as follows.
Short instructions: 1. Briefly depict the image. 2. Provide a concise overview of the presented image. 3. Summarize the visual elements in a succinct manner. 4. Give a clear, short explanation of the image. 5. Offer a compact interpretation of the provided image. 6. Share a brief account of the key features captured in the photo. 7. Relay a clear and concise description of the shown picture. 8. Render a succinct summary of the photo’s content. 9. Craft a compact narrative encapsulating the presented picture. 10. Create a brief, informative summary of the visual content.
Long instructions: 1. Elaborate on the specifics of the given image. 2. Offer an intricate explanation of the visual content. 3. Share a comprehensive rundown of the image presented. 4. Conduct a thorough analysis of the elements within the image. 5. Explain in detail the various aspects portrayed in the image. 6. Characterize the image through a well-detailed description. 7. Analyze the image comprehensively, delving into its details. 8. Illustrate the image through a descriptive explanation. 9. Examine the image closely and articulate its intricate details. 10. Craft an exhaustive depiction of the given image.
The image source of MMFundus dataset
We construct the large-scale multimodal fundus dataset (MMFundus) with images from 53 datasets. The data distribution is as follows: 1. HRF dataset [33]: 45 images, 2. INSPIRE-AVR dataset [34]: 40 images, 3. IOSTAR dataset [35]: 30 images, 4. RITE dataset [36]: 40 images, 5. G1020 dataset [37]: 1020 images, 6. GAMMA dataset [38]: 100 images, 7. ORIGA dataset [39]: 650 images, 8. REFUGE dataset [19]: 1200 images, 9. ODIR dataset [16]: 7000 images, 10. PALM dataset [40]: 1200 images, 11. RFMiD dataset [41]: 3200 images, 12. RFMiDv2 dataset [42]: 860 images, 13. APTOS dataset [17]: 3662 images, 14. DeepDRiD dataset [18]: 1600 images, 15. EyePACS dataset [43]: 35126 images, 16. IDRID dataset [44]: 516 images, 17. ADAM dataset [45]: 1200 images, 18. ACRIMA dataset [46]: 705 images, 19. MESSIDOR-2 dataset [47]: 1748 images, 20. JSIEC dataset [48]: 1000 images, 21. DeepEyeNet dataset [49]: 15708 images, 22. LAG dataset [50]: 4854 images, 23. PARAGUAY dataset [51]: 757 images, 24. PAPILA dataset [52]: 488 images, 25. STARE dataset [53]: 397 images, 26. FIVES dataset [54]: 800 images, 27. FUND dataset [55]: 179 images, 28. E-ophta dataset [56]: 463 images, 29. BRSET dataset [57]: 16266 images, 30. MuReD dataset [58]: 2208 images, 31. OIA-DDR dataset [59]: 12522 images, 32. SUSTech-SYSU dataset [60]: 1219 images, 33. Cataract dataset [61]: 601 images, 34. DGOCF dataset [62]: 9939 images, 35. BoVW dataset [63]: 2013 images, 36. HarvardGlaucoma dataset [64]: 1544 images, 37. RIM-ONE dataset [65]: 485 images, 38. CHAKSU dataset [66]: 1345 images, 39. DiaRetDB dataset [67]: 89 images, 40. LSD dataset [68]: 175 images, 41. GNG dataset [69]: 400 images, 42. AOD dataset [70]: 14813 images, 43. DHRF dataset [71]: 2757 images, 44. VietAI dataset [72]: 3435 images, 45. ToxoFundus dataset [73]: 411 images, 46. Papilledema dataset [74]: 1369 images, 47. BEH dataset [75]: 634 images, 48. WilliamHoyt dataset [76]: 856 images, 49. ROI dataset [77]: 1120 images, 50. ROD dataset [78]: 281 images, 51. BiDR dataset [79]: 2838 images, 52. AIROGS dataset [80]: 101442 images, 53. Private dataset: 33029 images.
Data Automatic Generation in MMFundus Dataset
Text Description Automatic Generation: We automatically generate text descriptions of the fundus images based on the original label information and generate corresponding sign labels for each image. The text description we design consists of three parts: normal/abnormal, specific diseases or conditions, and clinical explanations related to the disease. For example "Abnormal, Severe Diabetic Macular Edema [specific disease or condition 1], lots of hard exudates near to macula center observed [clinical explanation 1], Moderate Non-Proliferative Diabetic Retinopathy [specific disease or condition 2], Retinal hemorrhages or hard exudates observed [clinical explanation 2]." We can directly obtain information from the original label for normal/abnormal and specific diseases or conditions. In addition, we have designed a series of rules based on specific diseases or conditions to help generate corresponding clinical explanations, as shown below:
1. Abnormal, Mild Non-Proliferative Diabetic Retinopathy, Only microaneurysms observed. 2. Abnormal, Moderate Non-Proliferative Diabetic Retinopathy, Retinal hemorrhages or hard exudates observed. 3. Abnormal, Severe Non-Proliferative Diabetic Retinopathy, Many intraretinal hemorrhages or definite venous beading observed. 4. Abnormal, Proliferative Diabetic Retinopathy, Neovascularization or vitreous/preretinal hemorrhage. 5. Glaucoma, Abnormal optic disk color and unclear optic disk boundaries. 6. Abnormal, Cataract, Opacification of crystalline lens observed, Abnormal fundus color. 7. Abnormal, Hypertensive Retinopathy, Abnormal arterial vein ratio, Abnormal fundus color. 8. Abnormal, Myopia, Leopard fundus observed, Abnormal fundus color. 9. Abnormal, Media Haze, Opacity of media observed, Abnormal fundus color. 10. Abnormal, Drusen, Yellow or white extracellular deposits located between the retinal pigment epithelium (RPE) and Bruch’s membrane, Abnormal fundus color. 11. Abnormal, Branch Retinal Vein Occlusion, Occlusion of the central retinal vein, Abnormal fundus color. 12. Abnormal, Tessellation, The choroidal vessels are visible due to the reduced density of the pigments, Abnormal fundus color. 13. Abnormal, Epiretinal Membrane, A thin fibrous or cellular membrane that forms on the inner surface of the retina, Abnormal fundus color. 14. Abnormal, Laser Scars, Circular or irregular shaped scars on the retinal surface observed, Abnormal fundus color. 15. Abnormal, Macular Scar, Scar on the macula observed, Abnormal fundus color. 16. Abnormal, Central Serous Retinopathy, Fluid accumulation under the retina observed, Abnormal fundus color. 17. Abnormal, Optic Disk Cupping, The thinning of neuroretinal rim such that optic disc appears excavated, Abnormal fundus color. 18. Abnormal, Central Retinal Vein Occlusion, Occlusion of the central retinal vein, The presence of flame-shaped hemorrhages, Abnormal fundus color. 19. Abnormal, Tortuous Vessels, Marked tortuosity of the retinal blood vessels, Abnormal fundus color. 20. Abnormal, Asteroid Hyalosis, Numerous astroid bodies are dispersed in vitreous, Abnormal fundus color. 21. Abnormal, Optic Disc Pallor, Pale yellow discoloration of the optic disc, as well as absence of many small vessels, Abnormal fundus color. 22. Abnormal, Optic Disc Edema, Swelling of the optic disc. 23. Abnormal, Optociliary Shunt, Presence of prepapillary vascular loops or optociliary shunt vessels. 24. Abnormal, Anterior Ischemic Optic Neuropathy, Optic disc swelling and pallor. 25. Abnormal, Parafoveal Telangiectasia, Yellow, lipid-rich exudation or parafoveal graying or tortuous blood vessels. 26. Abnormal, Retinal Traction, Presence of traction and retinal traction detachment. 27. Abnormal, Retinitis, Presence of vitreous inflammation or intraretinal hemorrhage. 28. Abnormal, Chorioretinitis, The hard exudates observed. 29. Abnormal, Exudation, Retinal detachment. 30. Abnormal, Retinal Pigment Epithelium Changes, The structural changes in the RPE. 31. Abnormal, Macular Hole, A small retinal break located in the center of the fovea observed. 32. Abnormal, Retinitis Pigmentosa, The presence of bone-spicule deposits and arterial narrowing. 33. Abnormal, Cotton Wool Spots, The presence of soft exudates. 34. Abnormal, Coloboma, The missing of portion of tissue in both the choroid and retina. 35. Abnormal, Optic Disc Pit Maculopathy, The presence of optic disc pit. 36. Abnormal, Preretinal Hemorrhage, Boat-shaped hemorrhage which obscures the underlying retina. 37. Abnormal, Myelinated Nerve Fibers, Gray-white opaque lesions with feathery edges observed. 38. Abnormal, Hemorrhagic Retinopathy, The presence of flame-shaped hemorrhages. 39. Abnormal, Central Retinal Artery Occlusion, The presence of pale, whitening, and retinal swelling. 40. Abnormal, Tilted Disk, The tilting presence of the oval optic disc. 41. Abnormal, Cystoid Macular Edema, The presence of multiple cystoid areas in the macula and causes retinal edema. 42. Abnormal, Post-traumatic Choroidal Rupture, The breaks in the choroid, Bruch’s membrane, and RPE. 43. Abnormal, Choroidal Folds, The presence of folds in the choroid. 44. Abnormal, Vitreous Hemorrhage, The presence of extravasated blood in one of the spaces created around the vitreous body. 45. Abnormal, Macroaneurysm, Fusiform or round dilation of the retinal arterioles which occur in the temporal retina observed. 46. Abnormal, Vasculitis, The presence of inflammation of retinal blood vessels. 47. Abnormal, Branch Retinal Artery Occlusion, The presence of acute retinal artery obstructions. 48. Abnormal, Plaque, The plaque is present in retina. 49. Abnormal, Hemorrhagic Pigment Epithelial Detachment, The presence of hemorrhage from the Bruch’s membrane. 50. Abnormal, Collateral, New retinal vessels developed within the framework of existing vessel network. 51. Abnormal, Choroidal Neovascularization, The presence of subretinal fluid. 52. Abnormal, Cysticercosis, The presence of retinal edema and hemorrhage. 53. Abnormal, Giant Retinal Tear, The presence of retinal detachment and circumferential full-thickness tears of the retina. 54. Abnormal, Macular Edema, The macula region exhibits radially oriented cystoid pockets. 55. Abnormal, Optic Neuritis, The presence of optic disc swelling. 56. Abnormal, Retinal Detachment, The retina detaches from the retinal pigment epithelium. 57. Abnormal, Retinal Holes, small tears in the retina observed. 58. Abnormal, Retinal Tears, small breaches in the retina observed. 59. Abnormal, Hypertensive Retinopathy, Retinal hemorrhages or hard exudates observed. 60. Abnormal, Idiopathic Intracranial Hypertension, The presence of papilledema. 61. Normal, Healthy, Normal optic disk color and clear optic disk boundaries, Normal Macular color, Normal fundus color, No apparent retinopathy.
Dialogue Automatic Generation: The InternLM [31] and GPT-4 [32] are used to generate corresponding three rounds of dialogue data based on text descriptions. The prompt we use is as follows: "You will be provided with information about fundus images, including whether they are abnormal, optional specific disease, and diagnostic criteria. The fundus information is as follows: [Keyword]. You are an ophthalmologist who hopes to provide high-quality medical answers for patients. You need to provide the questions as the patients asking for fundus. The questions should be asking for more details and reasons. And you also need to provide a more detailed answer as an ophthalmologist. The answer should be appropriate to the question and the answer should be less than 200 words. The number of questions should be three." The text description of each image is filled as the keyword of the above prompt. After data cleaning, we can obtain three rounds of dialogue data corresponding to each fundus image, which can be used as the dialogue part of the MMFundus dataset.
Diagnostic Continuity in Patient Interaction: To evaluate the proficiency of models in adhering to and interpreting instructions, we quantify the diagnostic relevance of the responses from each group. Diagnostic relevance encapsulates the degree to which the content of the questions and answers align accurately and the extent to which patient inquiries are addressed comprehensively. Therefore, we employ the diagnostic relevance metrics as a basis for our ranking assessments. The outcomes of the empirical investigation are depicted in Figure g1-g4. Notably, our analysis reveals that in comparison to VisionUnite, both GPT-4V and Gemini Pro frequently yield responses that lack significance and demonstrate a suboptimal grasp of the posed questions. We postulate that this phenomenon may be attributed to instances where keyword triggers elicit unduly cautious reactions from the large models, thereby hindering their ability to engage with queries constructively. This observation underscores a prevalent dependency of contemporary large vision-language models on specific problem prompts, in contrast to the more resilient performance exhibited by VisionUnite. The consistency of VisionUnite in handling variously phrased questions with similar underlying meanings, which we term as ’prompt robustness’, further emphasizes its superior adaptability and understanding in diverse diagnostic scenarios. As shown in Extended Data Figure 3, we separately calculate the diagnostic relevance between each model and junior ophthalmologist in both correct and incorrect predictions. The results indicate that whether the prediction is correct or incorrect, VisionUnite’s ability to understand problems and follow instructions is comparable to the junior ophthalmologist. In the case of incorrect predictions, the diagnostic relevance of VisionUnite can still reach 2.084, higher than the ophthalmologists’ 1.951. The GPT-4V manifestation is in third place, with a diagnostic relevance of 1.86. The Gemini Pro performs the worst, with a diagnostic relevance of only 1.376. When the prediction is correct, ophthalmologists are often able to provide better clinical explanations, with a diagnostic relevance of 3.262. VisionUnite performs second, with a diagnostic relevance of 3.178. Gemini Pro and GPT-4V are consistent with themselves in each round of diagnostic relevance analysis. VisionUnite and ophthalmologists perform the best in the first round of diagnosis, while their diagnostic relevances slightly decrease in subsequent diagnoses. We believe it is due to the second and third rounds of diagnosis mainly focusing on clinical explanations and medical opinions. The problem is relatively general, so the performance of Gemini Pro and GPT-4V can be similar to that of VisionUnite and doctors, and it also leads to a slight decrease in the diagnostic relevance between VisionUnite and doctors in the second and third rounds.
Multiple-choice Evaluation: We conduct the multiple-choice evaluation and the results are shown in Figure 11. Compared with fine-tuned LLaVA-Med [24], Our proposed VisionUnite has achieved performance advantages on all ten diseases including Age-Related Macular Degeneration (AMD), Cataract, Central Serous Retinopathy (CSR), Diabetic Retinopathy (DR), Glaucoma, Media Haze, Myopia, Retinitis, Macular Edema and Tessellation. The performance advantage underscores the robustness and effectiveness of VisionUnite in accurately diagnosing a wide range of ocular conditions. A key factor contributing to the superior performance is its ability to embed explicit knowledge from sign labels. These sign labels represent specific features and indicators associated with each disease, allowing the model to make more informed and accurate predictions. By leveraging this explicit knowledge embedding, VisionUnite enhances its diagnostic accuracy compared to models that do not incorporate such detailed information.

### v2

@@ IV-J Diagnostic Errors Analysis between Ophthalmologist and Large Vision-Language Models
Figure 4: The consistent interpretation analysis of visual and language features between large vision-language models and the junior ophthalmologist (Doctor). Red represents inaccurate or irrelevant descriptions. Blue represents the accurate descriptions. The expert (senior ophthalmologist) evaluation is also provided.
IV-K Diagnostic Analysis of VisionUnite Assisted Doctor
To further explore the clinical decision-support capabilities of VisionUnite, we assess its impact on the diagnostic performance of primary ophthalmologists. We specifically evaluate the improvements in diagnostic accuracy and efficiency when VisionUnite assists ophthalmologists in the initial round which includes 180 questions. Our analysis shows that VisionUnite facilitates a 26.98% reduction in diagnostic time and a 29.44% increase in diagnostic accuracy overall illustrated in Figure 3. Notably, the use of VisionUnite leads to a remarkable 50% improvement in diagnostic accuracy and a 33% reduction in diagnostic time in cases of diabetic retinopathy. Furthermore, we examine the impact of VisionUnite on diagnoses involving five distinct physical signs detailed in Figure 3 i8-i12. The assistance is most significant for fundus hemorrhage exudates (FHE) and optic disc cupping (OCD), where VisionUnite helps improve diagnostic accuracy by 39.02% and 35.49%, respectively, while also reducing diagnostic time by 35.63% and 24.25%. In traditional fundus lesion screening, the reliance on the subjective judgment of ophthalmologists based on fundus photography can lead to variability in diagnosis owing to human factors such as fatigue, variability in training, and personal interpretative skills. This subjectivity can contribute to higher risks of missed diagnoses and misdiagnoses. In contrast, VisionUnite can standardize the interpretation of fundus images by consistently applying the same criteria to analyze and interpret data across all instances. In diseases with multifaceted presentations, such as diabetic retinopathy, VisionUnite can analyze multiple aspects of the disease, providing a holistic view that aids in a more comprehensive assessment. Additionally, predictive diagnostics further assist doctors in developing and evaluating treatment strategies.
IV-L Consistent Interpretation of Visual and Language Features
We explore the coherent integration of visual and textual elements within large vision-language models, an endeavor to assess the alignment between the textual descriptions and the corresponding visual data. Our investigation addresses the potential for ’illusory discrepancies’ within the output of these sophisticated models. Specifically, we list the outputs of VisionUnite against those from other large vision-language models, with a keen focus on discerning any textual outputs that may depict features not present within the associated imagery. For instance, in cases where the diagnosis pertained to age-related macular degeneration, and the imagery exclusively showcases the presence of drusen, we should evaluate whether the textual narrative unjustifiably extended the diagnosis to encompass additional manifestations such as hemorrhaging and exudation as shown in Figure 4. Such a detailed examination is pivotal for understanding and mitigating the tendency of these models to over-interpret or misalign textual descriptions with their visual counterparts, thereby ensuring a more accurate and reliable diagnostic output. Meanwhile, we also present a series of evaluation criteria for the responses of the models and junior ophthalmologists in this section.
IV-M Ablation Studies of Proposed Components
To validate individual contribution of our proposed component, we conduct ablation studies across our multi-round VQA evaluation framework. Table XIV presents comprehensive results demonstrating the necessity and efficacy of each component. The results demonstrate that each proposed component contributes meaningfully to diagnostic performance. Individual component analysis demonstrates that the Vision Adapter with classification loss ($L_{CLS}$) achieves the highest standalone performance (75.95%), followed closely by the contrastive learning component ($L_{CLIP}$) at 75.14%, while the Vision Projector yields 74.63% when used in isolation. This performance hierarchy indicates that sign-level feature extraction and image-text alignment represent the most fundamental capabilities for ophthalmological diagnosis, with the Vision Adapter’s explicit clinical knowledge encoding providing slight advantages over general visual-linguistic alignment. The relatively lower performance of the Vision Projector in isolation suggests its primary role as a facilitating component that enhances the effectiveness of other modules rather than serving as a standalone diagnostic feature extractor.
TABLE XIV: The ablation study of each proposed component on multi-rounds VQA dataset using SBERT Similarity (%). Due to the dependence of $L_{CLS}$ on Vision Adapter, Vision Adapter and $L_{CLS}$ are considered as the same component.
$L_{CLIP}$ Vision Adapter($L_{CLS}$) Vision Projector Round 1 Round 2 Round 3 Overall $\uparrow$
76.83 75.04 73.55 75.14
77.62 75.37 74.86 75.95
76.41 74.28 73.19 74.63
82.45 77.81 76.93 79.06
79.35 76.79 75.91 77.35
81.42 77.15 76.34 78.30
83.46 78.53 77.82 79.94
IV-N Ablation Studies of Pretraining and/or Finetuning Datasets
To validate the individual contributions of pretraining and fine-tuning datasets and ensure fair comparison with baseline configurations, we conduct systematic ablation studies across multiple dataset combinations. This analysis isolates the performance impact of each dataset component, providing transparent evaluation of our methodological contributions. We evaluate six distinct configurations using our established multi-round VQA evaluation protocol with SBERT similarity assessment, ranging from a baseline configuration using only MMFundus fine-tuning to our complete configuration incorporating all pretraining datasets. Table XV presents comprehensive results demonstrating systematic performance improvements across dataset configurations, with the baseline configuration achieving 74.69% overall performance and establishing a clear comparative foundation. The results reveal that MS COCO pretraining contributes a 2.96% improvement (77.65% vs 74.69%), demonstrating the value of general vision-language representations for ophthalmological applications, while PMC-OA pretraining provides a 2.07% improvement (76.76% vs 74.69%), validating the importance of biomedical domain knowledge integration. The combined pretraining datasets yield a 4.69% improvement (79.38% vs 74.69%), indicating complementary benefits from diverse pretraining sources and confirming the synergistic effects of our comprehensive training approach. It should be noted that we specifically modified the format and style of RET. annotations to ensure that VisionUnite could generate responses in the same style without finetuning on MMFundus. The results demonstrate consistent performance patterns across all three dialogue rounds, with the complete configuration achieving optimal performance in Round 1 (83.46%), Round 2 (78.53%), and Round 3 (77.82%).
TABLE XV: The ablation study of pretraining and/or finetuning datasets on multi-rounds VQA dataset using SBERT Similarity (%). RET. indicates Retina Image Bank dataset.
Pretraining datasets Finetuning dataset Round 1 Round 2 Round 3 Overall $\uparrow$
MS COCO PMC-OA RET. MMFundus
76.35 74.58 73.14 74.69
80.92 76.44 75.59 77.65
79.27 76.05 74.96 76.76
82.78 78.12 77.24 79.38
75.51 73.83 72.33 73.89
83.46 78.53 77.82 79.94
V Discussion
The diagnostic efficacy of VisionUnite in ophthalmic conditions has been extensively validated in various clinical settings, indicating improvements in initial screenings, thereby enhancing healthcare efficiency, especially in under-resourced areas. The image analysis and diagnostic capabilities of VisionUnite are comparable to junior ophthalmologists, sometimes even surpassing them. It outperforms other large vision-language models like GPT-4V and Gemini Pro by providing more accurate diagnoses and clearer explanations, crucial for clinical use. VisionUnite also accelerates disease diagnosis, supports patient interaction, and aids in educational efforts for healthcare professionals by providing detailed reports on fundus photography, a feature that traditional models without textual descriptions lack. Our model employs a modular design specifically engineered for component-wise optimization, allowing selective fine-tuning of individual components based on computational resources. This approach enables a progressive deployment strategy where resource-constrained environments can implement a reduced version with fewer classification tasks while maintaining core diagnostic functionality. While preliminary evaluations suggest VisionUnite may approximate performance characteristics of junior ophthalmologists under controlled conditions, significant limitations constrain its clinical utility. VisionUnite demonstrates suboptimal performance in several ophthalmic conditions, primarily due to data scarcity within datasets. Currently restricted to fundus imaging, VisionUnite exhibits the narrow assessment capabilities typical of current AI diagnostic tools, with limited classification labels that restrict detailed diagnostic assessments. Future developments may expand to additional ophthalmic imaging modalities and implement more granular sign-level labeling to enhance diagnostic accuracy. As technology and data availability improve, VisionUnite is anticipated to handle a wider range of conditions, potentially identifying connections between ocular signs and systemic diseases, thus advancing both ophthalmic and general medical diagnostic processes.
VI Conclusion
In this study, we propose VisionUnite, which represents a significant advancement as a large vision-language foundation model for ophthalmology with clinical knowledge. Its distinguishing feature lies in its open diagnostic capability for eye diseases, eliminating the requirement for predefined disease ranges and aligning more closely with the demands of clinical diagnosis. Furthermore, the utilization of VisionUnite holds the promise of expediting the identification of previously undiscovered connections between diseases and ocular features, contributing to the refinement of diagnosis.
Acknowledgments
This research is supported by Shanghai Artificial Intelligence Laboratory. This work is partially supported by the National Key R&D Program of China (NO.2022ZD0160102, NO.2022YFC2502800), partially supported by the National Natural Science Foundation of China (62272450, 82471066), and partially supported by the China Postdoctoral Science Foundation (Grant No.2022M721743). This work was done during Zihan Li’s internship at Shanghai Artificial Intelligence Laboratory.